# Causality for Earth Science - A Review on Time-series and Spatiotemporal Causality Methods

### v1

@@ Abstract
This survey paper covers the breadth and depth of time-series and spatiotemporal causality methods, and their applications in Earth Science. More specifically, the paper presents an overview of causal discovery and causal inference, explains the underlying causal assumptions, and enlists evaluation techniques and key terminologies of the domain area. The paper elicits the various state-of-the-art methods introduced for time-series and spatiotemporal causal analysis along with their strengths and limitations. The paper further describes the existing applications of several methods for answering specific Earth Science questions such as extreme weather events, sea level rise, teleconnections etc. This survey paper can serve as a primer for Data Science researchers interested in data-driven causal study as we share a list of resources, such as Earth Science datasets (synthetic, simulated and observational data) and open source tools for causal analysis. It will equally benefit the Earth Science community interested in taking an AI-driven approach to study the causality of different dynamic and thermodynamic processes as we present the open challenges and opportunities in performing causality-based Earth Science study.
@@ 1 Introduction
Earth science is the study of the Earth’s physical properties, structure, and processes. It encompasses a wide range of topics including geology, meteorology, oceanography, and environmental science (Smith and Pun, 2013). Earth science research relies purely on climate models that replicate complex dynamical systems. Climate models allow scientists to better understand how the Earth’s climate system works and how it might change in the future. These models can be used to make predictions about future climate change and to test different scenarios, for instance studying the global carbon footprint to reduce greenhouse gas emissions. However, climate models are computationally expensive and require a lot of computing power to run. This is because they need to simulate many different dynamic processes that occur in the atmosphere, oceans, and on land. For instance a simulation that works on cloud formation or generating ocean currents.
@@
In recent decades, there has been a significant increase in the availability of large-scale climate data from various observational sources (such as satellite remote sensing, station-based or field site measurements) and Earth system model outputs (Guo et al., 2015). This data, along with the increasing computational power, has opened up new ways to use data-driven methods for observational causal discoveries without relying on the correlation and trend analyses of the data (Rubin, 2005). In Earth system sciences, causal structure discovery (CSD) models are making inroads into several domains (Melkas et al., 2021). Climate causal studies pose the changing trend of applying the current state-of-the-art climate data not only to correlation and regression methods but also to causal inference methods. For instance, climate researchers have realized that climate simulations introduce ambiguous values to the datasets which makes them not applicable in decision-making applications (Ebert-Uphoff and Deng, 2012). After understanding the importance of data-driven methods in climate science, scientists have started to explore different causal structure algorithms and Bayesian networks to get deeper insights into causal hypotheses and the evaluation of physical models (Ebert-Uphoff and Deng, 2012). In the past, Bayesian models were used to compute the probability of climate data where it is needed for forecasting and risk assessment tools. Now, Bayesian models are significantly used in climate science to test the causal hypothesis to develop decision-making algorithms. A recent research study (Nowack et al., 2020) evaluated the ability of the Coupled Model Intercomparison Project Phase 5 (CMIP5) to simulate atmospheric dynamical interactions using the causal model evaluation (CME) framework. This study highlighted the potential of the CME framework to offer a pathway to reducing uncertainties in climate change projections by quantifying differences between models and observations. To sum it up, causality methods can play a crucial role in Earth Science studies by helping researchers understand and quantify the cause-and-effect relationships within complex Earth systems for a variety of usecases. Some of these include: identifying the primary drivers of environmental changes, attributing specific climate events like volcanic eruptions, identifying causes of extreme weather events such as hurricanes, droughts and floods, helping in the prediction of natural hazards such as earthquakes and tsunamis, designing effective pollution control policies, understanding El Niño/La Niña events and their global climate impacts, etc.
@@
This survey paper is therefore an effort to provide an overview of the current state of knowledge on causality in Earth Science. It will serve as a primer to help understand key concepts in causality and methods that are commonly used. In addition to existing survey papers (Table 1), our survey paper will help identify gaps in current causality-based Earth Science research and highlight areas where causal methods could be used to improve our understanding of the Earth’s system. The paper is organized as follows. Section 2 enlists some open challenges in causality-based study of Earth Science problems. Section 3 explains the concept of causal structure learning or ”Causal Discovery”. Here, we first explain the foundational concepts of causal discovery such as key terms, causal assumptions and the relevant evaluation metrics for the structure learning task. We then enlist several approaches to perform causal discovery on time-series and spatiotemporal data. Finally, we share the applications of causal discovery in Earth Science as found in the literature. Section 4 explains the process of estimating causal effects, also known as ”Causal Inference”. Just like the previous section, we first explain the key terms, causal assumption and evaluation metrics related to Causal inference. The subsequent topics elicits the various state-of-the-art methods introduced for time-series and spatiotemporal causal inference along with their strengths and limitations. This section further describes the existing applications of several causal inference methods for answering specific Earth Science questions such as extreme weather events, sea level rise, etc. In Section 5, we share a list of resources including Earth Science datasets (synthetic, simulated and observational data) and open source toolboxes for causal analysis. Lastly, we conclude our work and share potential research pathways in Section 6. We hope our paper equally benefits both the Data Science community interested in data-driven causal study and the Earth Science community interested in taking an AI-driven approach to study the causality of different dynamic and thermodynamic processes.
@@
What factors drive the loss of glacier mass in both the Arctic and Antarctic regions, and what implications does
this have for global weather patterns?
@@
How does sea level rise impact water-related extreme events (e.g., floods and droughts)?
@@ 2 Open Challenges in Causality-based Earth Science Study
Due to the multi-way interactions between the atmosphere, hydrosphere, cryosphere, geosphere, and biosphere, studying causality between them is a challenging but important task, which makes it an area of high interest within the Earth Science community. Here, we have identified several fundamental questions within the Earth Science domain, particularly concerning the rapidly increasing emissions of greenhouse gases and the consequential impacts of climate change, which may benefit from causality methods. These inquiries pose significant challenges due to several factors: 1) Data availability may be inadequate or insufficient, hindering comprehensive analysis. 2) Climate data is inherently complex, with multiple dimensions, making analysis and interpretation challenging. 3) The complex interaction among numerous variables complicates the disentanglement of causal relationships and feedback mechanisms. 4) Confounding effects further hinder the clarification of causality. 5) The inability to conduct controlled laboratory experiments in climate science limits our understanding of causal relationships. 6) The current state of climate models may lack the sophistication and precision necessary to accurately capture the complex interactions among variables. In Table 2, we enlist some of the causality-related open questions in different Earth Science domains (Masson-Delmotte et al., 2021; Pörtner et al., 2022). In Table 3, we identify some of the challenges in performing observational causal discovery and causal inference in various domains of Earth Science and the opportunities they pave for researchers.
@@ 3 Causal Discovery
Causal discovery refers to the process of discovering causal relations and uncovering underlying data generation process in real world data, helping researchers make more informed decisions, predict outcomes, and understand how changes in one variable, state or process may lead to changes in others (Pearl, 1988). In recent years, several graph-based algorithms are proposed for discovering the causal structures from empirical data (Spirtes et al., 2000; Tian and Pearl, 2013). The important causal connections are discovered purely based on the statistical analyses of observation data. This section presents an overview of time-series and spatiotemporal causal discovery, the key concepts in causality, and common approaches for performing causal structure learning for both temporal and spatial datasets. It further includes brief overview of renowned causal discovery methods, their limitations and applications in Earth Science study. To begin with, we have enlisted the key terminologies in causal discovery in Table  7.
@@
Causal discovery models It aims to discover causal structures from data
Causal inference Estimating the strength of the causal relationship between two variables
@@
Probabilistic Graphical Model It captures conditional independence relationships between interacting random variables.
Bayesian Network Model A probabilistic graphical model that represents the conditional dependencies of a set of
random variables via a directed acyclic graph (DAG)
Latent parameter Unknown / hidden parameters or hypotheses
Conditional independence Nodes that are not causally related / connected
Conditional dependence Nodes that are causally related / connected
Markov Network An undirected graphical model with a set of random variables having a Markov property
@@
Markov equivalence class A set of DAGs that shares the same set of conditional independencies
Mutual Information Network A measure of mutual information between two variables in the Bayesian network
@@ 3.1 Causal Assumptions
There are certain foundational principles and conditions, referred to as causal assumptions, that underlie the process of inferring causal relationships between variables. Different causal discovery methods may have different sets of assumptions. Some of the common causal assumptions are given below:
@@
If for every measured variable in the data, all of its common causes are also measured, then the causal sufficiency assumption holds. As per this condition, a pair of variables is causally sufficient if all the common causes are measured, i.e. there are no latent confounders (Nogueira et al., 2021). Causal sufficiency is a very strong assumption as there is no such thing as a closed world. In real-world settings, it is not always the case that all possible causes are measured and hence, this assumption is often violated (Pellet and Elisseeff, 2008). Causal insufficiency occurs when for a set of variables V, there exists variables not present in V that are direct causes of more than one variable in V. In Figure 2, the variables $A,C,B$ are causally insufficient since another variable D outside of set V is the common cause of B & C.
@@
The causal Markov assumption states that every node in a causal DAG G is independent of its non-descendants when conditioned on its parents. The idea behind the CMC is that when a variable’s effects (descendents) are ignored, all the relevant probabilistic information about that variable can be obtained from its direct causes (parents) (Scheines, 1997). CMC can also be explained in terms of d-separation. In a causal DAG G, all variables that are d-separated will be conditionally independent of each other in the respective probability distribution (Weinberger, 2018). This also implies that, among three disjoint nodes, one node must block all paths between the other two nodes (Nogueira et al., 2021). The causal Markov assumption is true in a DAG which is causally sufficient i.e. have no latent confounders.
Figure 3: An example to demonstrate the causal Markov property.
The CMC for the variables in the causal graph of Figure 3 is as follows:
C $\perp$ B |||| A (C is independent of its non-descendant B conditioned on its parent A)
B $\perp$ C |||| A (B is independent of its non-descendant C conditioned on its parent A)
@@
The causal faithfulness assumption can be stated as follows. In a causal DAG G, no conditional independence between the variables holds unless entailed by the causal Markov condition (CMC). That is, CFC specifies which variables in a DAG will be probabilistically dependent (Weinberger, 2018). This is the converse of the Markov condition since CMC specifies which variables will be conditionally independent given their parents. With respect to d-separation, CFC means that except for the d-separated variables in a DAG, all other variables are dependent. In Figure 3, variables (A,C) and (A,B) are dependent. Based on CFC, it can be interpreted that no observed interdependency in the data is accidental, but results from the underlying causal mechanism (Druzdzel, 2009).
@@
The acyclicity assumption states that no feedback loops are allowed in a causal graph. That is, there can be no directed paths from a variable back to itself similar to the structure of a directed acyclic graph (DAG). This is a very basic property of a causal graph Figure 4b) that must be fulfilled to represent the causal relationships accurately. Therefore, a causal graph is a DAG where no variable is its ancestor or its descendant (Greenland and Pearl, 2007).
@@
Causal discovery approaches heavily rely on the available data and their assumptions. The data for performing any causal analysis can be either observational or interventional in nature, or both. To conduct causal discovery, often different types of assumptions about the data distribution are considered. They include data being linear or nonlinear, continuously valued or discrete valued, stationary or non-stationary (for temporal data), etc. Often assumptions about the varied noise distributions such as Gaussian, Gumbel, or Exponential noise to which the data may belong are made. Apart from these, some common assumptions about data are the existence of any sampling or selection bias, missing data, latent variables or noise, etc.
@@ 3.2 Evaluation Metrics
True Positive Rate (TPR):
TPR is the probability that an actual positive will test positive. Assuming the threshold $t$ of probability of an edge $p(a_{ij})$ ranges from $(0,1)$, TPR is defined as
$TPR_{t}=|\frac{\{(i,j):p(a_{ij})\geq t\}\cap S}{S}|$ (1)
where $S$ is the set of ground truth edges (i.e., ${(i,j):a^{*}_{ij}=1})$.
False Positive Rate:
False positive rate (FPR) is the probability that an actual negative will test positive. Assuming the threshold $t$ of probability of an edge $p(a_{ij})$ ranges from $(0,1)$, $FPR$ is defined as
$FPR_{t}=|\frac{\{(i,j):p(a_{ij})\geq t\}\cap\hat{S}}{\hat{S}}|$ (2)
where $\hat{S}$ is the set of ground truth missing edges (i.e., ${(i,j):a^{*}_{ij}=0}$).
Structural Hamming Distance (SHD):
Structural Hamming Distance (SHD) is the number of operators required to make two PDAGs match. The operators include adding or removing an undirected edge, and adding, removing or reversing the orientation of an edge. Let $\mathbb{P}$ be the space of PDAGs over $p$ variables. SHD is defined as:
$\begin{split}SHD:&\mathbb{P}\times\mathbb{P}\Rightarrow\mathbb{N}\\ &(\mathcal{G,H})\mapsto{(i,j)\in\mathbf{V^{2}}}\end{split}$ (3)
Structural Intervention Distance (SID):
Structural intervention distance (SID) is the number of vertice pairs $(i,j)$ for which the estimate DAG correctly predicts intervention distributions within the class of distributions that is Markov with respect to another DAG.
Area Over Curve (AOC):
Area over curve (AOC) simply area under curve (AUC) with $1-AUC$. The curve is plotted with $TPR_{t}$ against $FPR_{t}$ as $(FPR_{t},TPR_{t})$, where $t\in[0,1]$.

### v2

@@ Abstract
This survey paper covers the breadth and depth of time-series and spatiotemporal causality methods, and their applications in Earth science. The paper first introduces the concepts of causal discovery and causal inference, followed by the underlying causal assumptions, evaluation techniques and key terminologies of the two domain areas. The paper elicits the various state-of-the-art methods introduced for time-series and spatiotemporal causal analysis along with their strengths and limitations. The paper further describes the existing applications of several methods for answering specific Earth science questions such as extreme weather events, sea level rise, teleconnections, etc. Our survey paper will benefit the Earth science community interested in taking an AI-driven approach to study the causality of different dynamic and thermodynamic processes as we present the open challenges and opportunities in performing causality-based Earth science study. It will also serve as a primer for data science researchers interested in data-driven causal study as we share a holistic list of resources, such as Earth science datasets (synthetic, simulated and observational data) and open source tools for causal analysis.
@@ 1 Introduction
Earth science is a broad field that studies the Earth’s composition, processes, and physical characteristics (Smith and Pun, 2013). It encompasses several sub-disciplines like geology (rocks and minerals), meteorology (atmosphere and weather), oceanography (oceans and their properties), and environmental science (interactions between living things and the environment). Climate models, complex computer simulations, play a vital role in Earth science research. These models enable predictions about climate trends and allow for the testing of various scenarios, such as analyzing the global carbon footprint in order to mitigate greenhouse gas emissions. Despite their importance, climate models are computationally demanding and necessitate significant computing resources. This requirement stems from the need to simulate numerous dynamic processes occurring in the atmosphere, oceans, and on land, such as cloud formation and ocean current generation.
@@
In recent decades, there has been a significant increase in the availability of large-scale climate data from various observational sources (such as satellite remote sensing, ground-based, sea-borne, or air-borne systems) and climate model outputs (Guo et al., 2015). This large volume of data has opened up new ways to use data-driven methods for observational causal discoveries without relying on the correlation and trend analyses of the data (Rubin, 2005). Causal structure discovery (CSD) models are becoming increasingly valuable across various research areas in Earth system sciences (Melkas et al., 2021). Climate causal studies pose the changing trend of applying the current state-of-the-art climate data not only to correlation and regression methods but also to causal inference methods. For instance, climate researchers have realized that climate simulations introduce ambiguous values to the datasets which makes them not applicable in decision-making applications (Ebert-Uphoff and Deng, 2012). After understanding the importance of data-driven methods in climate science, scientists have started to explore different causal structure algorithms and Bayesian networks to get deeper insights into causal hypotheses and the evaluation of physical models (Ebert-Uphoff and Deng, 2012). To sum it up, causality methods can play a crucial role in Earth science by helping researchers understand and quantify the cause-and-effect relationships within complex Earth systems for a variety of usecases (Nowack et al., 2020). Some of these include: identifying the primary drivers of environmental changes, attributing specific climate events like volcanic eruptions, identifying causes of extreme weather events such as hurricanes, droughts and floods, helping in the prediction of natural hazards such as earthquakes and tsunamis, designing effective pollution control policies, understanding El Niño/La Niña events and their global climate impacts, etc.
@@
This survey paper presents a comprehensive overview of the current understanding of causality in Earth science. It serves as an introduction to key concepts in causality and the commonly used methods in this field. Additionally, this survey identifies gaps in current research and highlights areas where causal methods could improve our understanding of the Earth’s system. The paper is organized as follows: Section 2 discusses the open challenges in causality-based studies of Earth science problems. Section 3 explores the concept of causal structure learning, or ”Causal Discovery”. This section covers foundational concepts such as key terms, causal assumptions, and relevant evaluation metrics for structure learning. It also details various approaches for performing causal discovery on time-series and spatiotemporal data and reviews their applications in Earth science. Section 4 delves into ”Causal Inference” techniques for estimating causal effects. It also covers state-of-the-art methods for time-series and spatiotemporal causal inference, detailing their strengths and limitations and providing an example of their application in Earth science. Section 5 lists resources for synthetic, simulated, and observational data commonly used in Earth science, as well as open-source toolboxes for causal analysis. Section 6 wraps up the paper with a discussion on potential research directions. We aim for this paper to benefit both the Data Science community, which is interested in data-driven causal studies, and the Earth science community, which seeks to use AI-driven approaches to explore the causality of various dynamic and thermodynamic processes.
@@
What factors drive the loss of glacier mass in both the Arctic and Antarctic regions, and how does it
contribute to global weather patterns?
@@
How does sea level rise impact extreme events such as, floods and droughts?
@@ 2 Open Challenges in Causality-based Earth science Study
Due to the multi-way interactions between the atmosphere, hydrosphere, cryosphere, geosphere, and biosphere, studying causality between them is a challenging but important task. Here, we have identified several fundamental questions within the Earth science domain, particularly concerning the rapidly increasing emissions of greenhouse gases and the consequential impacts of climate change, which may benefit from causality methods. These inquiries pose significant challenges due to several factors: 1) Data availability may be inadequate or insufficient, hindering comprehensive analysis. 2) Climate data is inherently complex, with multiple dimensions, making analysis and interpretation challenging. 3) The complex interaction among numerous variables complicates the disentanglement of causal relationships and feedback mechanisms. 4) Confounding effects further hinder the clarification of causality. 5) The inability to conduct controlled laboratory experiments in climate science limits our understanding of causal relationships. 6) The current state of climate models may lack the sophistication and precision necessary to accurately capture the complex interactions among variables. In Table 2, we enlist some of the causality-related open questions in different Earth science domains (Masson-Delmotte et al., 2021; Pörtner et al., 2022). In Table 3, we identify some of the challenges in performing observational causal discovery and causal inference in various domains of Earth science and the opportunities they pave for researchers.
@@ 3 Causal Discovery
Causal discovery refers to the process of discovering causal relations and uncovering underlying data generation process in real world data, helping researchers make more informed decisions, predict outcomes, and understand how changes in one variable, state or process may lead to changes in others (Pearl, 1988). In recent years, several graph-based algorithms are proposed for discovering the causal structures from empirical data (Spirtes et al., 2000; Tian and Pearl, 2013). The important causal connections are discovered purely based on the statistical analyses of observation data. This section presents an overview of time-series and spatiotemporal causal discovery, the key concepts in causality, and common approaches for performing causal structure learning for both temporal and spatial datasets. It further includes brief overview of renowned causal discovery methods, their limitations and applications in Earth science study. To begin with, we have enlisted the key terminologies in causal discovery in Table  4.
@@
Causal discovery models Discovers causal relationships between variables
@@
Probabilistic Graphical Model Captures conditional independence relationships between random variables.
Bayesian Network Uses probability theory and a Directed Acyclic Graph (DAG) to quantify the relationships between variables.
Latent parameter Unknown / hidden parameters
Conditional independence Nodes that are not causally related
Conditional dependence Nodes that are causally related
Markov Equivalence Class A set of Directed Acyclic Graphs (DAGs) that may vary in structure (with different node arrangements and arrow directions) but all represent the same set of conditional independence relationships among the variables.
@@ 3.1 Causal Assumptions
There are several foundational principles and conditions known as causal assumptions that are essential for inferring causal relationships between variables. Various causal discovery methods rely on different sets of these assumptions. Below are some commonly recognized causal assumptions:
@@
According to this condition, a pair of variables is considered causally sufficient if all their common causes are measured, meaning there are no latent confounders (Nogueira et al., 2021). Causal sufficiency is a very stringent assumption because the concept of a closed world does not exist. In real-world scenarios, it is often the case that not all possible causes are measured, leading to frequent violations of this assumption (Pellet and Elisseeff, 2008). Causal insufficiency occurs when, for a set of variables $V$, there are variables not included in $V$ that are direct causes of more than one variable within $V$. In Figure 2, the variables $A,C,B$ are causally insufficient because another variable $D$ outside of set $V$ is the common cause of $B$ and $C$.
@@
The principle behind the CMC is that all pertinent probabilistic information about a variable can be derived from its direct causes (parents), even if the variable’s descendants (effects) are ignored (Scheines, 1997). The CMC can also be understood through d-separation. In a causal DAG $G$, variables that are d-separated will be conditionally independent in the corresponding probability distribution (Weinberger, 2018). This implies that among three disjoint nodes, one node must block all paths between the other two nodes (Nogueira et al., 2021). The causal Markov assumption holds in a DAG that is causally sufficient, meaning there are no latent confounders. The CMC for the variables in the causal graph of Figure 3 is as follows:
B $\perp$ C |||| A (B and C are independent of each other conditioned on their common parent A)
Figure 3: An example to demonstrate the causal Markov condition/property.
@@
The causal faithfulness assumption can be described as follows: In a causal DAG $G$, no conditional independence between variables exists unless it is implied by the causal Markov condition. Essentially, CFC identifies which variables in a DAG will be probabilistically dependent (Weinberger, 2018). This is the opposite of the Markov condition, which specifies which variables will be conditionally independent given their parents. In terms of d-separation, CFC indicates that except for the d-separated variables in a DAG, all other variables are dependent. In Figure 3, the variables $(A,C)$ and $(A,B)$ are dependent. According to CFC, this implies that no observed interdependency in the data is accidental, but rather results from the underlying causal mechanism (Druzdzel, 2009).
@@
The acyclicity assumption asserts that no feedback loops are permitted in a directed causal graph. This means there cannot be a directed path from a variable leading back to itself. This is a fundamental property of a causal graph, as shown in Figure 4, which must be satisfied to accurately represent causal relationships. Therefore, a causal graph is often represented using a DAG where no variable is an ancestor or descendant of itself (Greenland and Pearl, 2007). Adhering to the acyclicity condition is essential for the proper application of many causal inference techniques, as it prevents the complications that arise from circular causality and helps maintain the integrity of the causal model.
@@
Causal discovery approaches significantly depend on the available data and the assumptions made about it. The data used for causal analysis can be observational, interventional, or both. To perform causal discovery, various assumptions about the data distribution are often considered. These assumptions include whether the data is linear or nonlinear, continuously valued or discrete valued, and whether it is stationary or non-stationary (for temporal data). Additionally, assumptions about different noise distributions, such as Gaussian, Exponential, or Gumbel noise, to which the data may belong, are commonly made. Other frequent assumptions about the data include the presence of any sampling or selection bias, missing data, latent variables, or noise. These assumptions are crucial because they provide the necessary framework for accurately inferring causal relationships from data. Without these assumptions, the causal inferences drawn could be misleading or incorrect, ultimately compromising the conclusions and applications derived from the causal analysis.
@@ 3.2 Evaluation Metrics
True Positive Rate (TPR)
is the likelihood ratio of true causal edges to be identified correctly by the causal method. Assuming the threshold $t$ of the probability of an edge $Prob(e_{ij})\in(0,1)$, the calculation of TPR is:
$TPR_{t}=|\frac{\{(i,j):Prob(e_{ij})\geq t\}\cap S}{S}|$ (1)
where $S$ is the set of edges whose ground truth is positive.
False Positive Rate (FPR)
is the likelihood ratio of absence of causal edges to be identified incorrectly by the causal method. Assuming the threshold $t$ of the probability of an edge $Prob(e_{ij})\in(0,1)$, $FPR$ is defined as
$FPR_{t}=|\frac{\{(i,j):Prob(e_{ij})\geq t\}\cap\hat{S}}{\hat{S}}|$ (2)
where $\hat{S}$ is the set of edges that do not exist in the true causal graph.
Structural Hamming Distance (SHD)
represents number of operators required to make two partial DAGs match. The operators include adding or removing an undirected edge, and adding, removing or reversing the direction of an edge. The algorithm of SHD can be found in(Tsamardinos et al., 2006).
Structural Intervention Distance (SID)
measures number of vertex pairs $(i,j)$ in the estimated Directed Acyclic Graph (DAG) that correctly predict the intervention distributions among the distributions that are Markov equivalent with respect to another DAG.
Area Over Curve (AOC)
simplifies area under curve (AUC) with $1-AUC$. The curve is plotted with $TPR_{t}$ against $FPR_{t}$ as $(FPR_{t},TPR_{t})$, where $t\in[0,1]$.